Shift summary for owl 7/12/99

S. Elizabeth Turner (eturner@kudzu)
Mon, 12 Jul 1999 07:04:19 -0500

On Shift: RK, SET

Hours of beam available at all intensities: 7
Hours of beam available at nominal intensity: 7
Hours of E832 data (special runs not included): 5.75
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: tape change (15 minutes) and
pipeline problems (~1 hour)

========================================================================
Runs ended this shift: 13860 BEAM_832

Tapes brought to FCC this shift: none
========================================================================

Accesses during shift: none
========================================================================

Overview of shift:
Regular E832 data taking, interupted by pipeline problems.
Rebooting a few pipeline crates seemed to fix the problem,
but not before losing an hour of beam.

Some scaler alarms were added, including DAQDEAD_SEC, HA2/REG
and L2/L1 ratio. SEM alarm set to 7.1E12 !!!

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13860 1:00 DAQDEAD_SEC scaler revived! But it counts 2 seconds
per spill because DYC 53 BUSY was plugged into the
same BUSY group as the BAFs and DDD vetos.

285 1:30 Quad drops out of laser illum. Reset
pipe, hcc, etot -- none of these fix
the problem.
2:00 End run, change tapes.
Note that spills 285-end are no good.

13861 2:15 Drt fresh start. Start run. Laser illum ok
now (was fixed by drt_fresh_start & prayer).

Moved DYC 53 BUSY and FULL to spare inputs in the
nominal DYC BUSY and FULL groups (tested with NIM
level inputs too). Now DAQDEAD_SEC counts 0.01 sec
per spill. SET SCALER ALARM to 0.5 seconds to catch
daq dead time problems. HA scaler alarm added too.

In the flurry of the above excitement we forgot
to change the magnet polarity. We skipped a muon
run for now also.

7 2:31 ktevana errors; then pipeline erorrs.
Had to reboot a few crates; lost spills 7-18.

41-42 3:16 dyc53 full -- kfrend fixed it, lost these spills.

67-68 3:52 dyc53 full -- kfrend, lost these spills.

148 5:40 Called MCR -- asked them to boost intensity
from 6.2E12 to request of 6.8E12.

177 6:20 Called MCR -- asked them to lower intensity
(we've been getting as high as 7.3E12 which
causes too much deadtime)

195 6:40 SEM looks normal (6.8E12) but all scalers are
counting 3 orders of magnitude below their
normal values (just for this one spill) -- don't
know what happened.

========================================================================
Follow up Notes on Previous Shift

Keep an eye on the +/- 24V supply to RR160.
--button stayed red for the whole shift but voltages
were never very far out of bounds so we didn't worry
too much.

========================================================================
Items to Follow up on Next Shift

MUON RUN and change magnet polarity